### v1

@@ Abstract
In this paper, we contribute a formal verification of the main construction required to establish said universal pair. In doing so, we markedly extend the Isabelle AFP entry on multivariate polynomials [17], formalize parts of a number theory textbook [15], and develop classical theory on Diophantine equations [14] in Isabelle. Additionally, our work includes metaprogramming infrastructure designed to efficiently handle complex definitions of multivariate polynomials. Our mathematical draft has been formalized while the mathematical research was ongoing, and benefitted largely from the help of the theorem prover. We reflect how the close collaboration between mathematician and computer is an uncommon but promising modus operandi.
@@ 1.1 Hilbert’s Tenth Problem
The tenth problem on David Hilbert’s 1900 list of twenty-three mathematical problems asks for an algorithm to decide the solvability of any arbitrary Diophantine equation [7]. Like many problems on his list, Hilbert’s Tenth Problem has provided inspiration for much mathematical research in the last century, and still does so until today. While the original problem was solved over fifty years ago [10], this paper presents the formal verification of results related to a new quantitative theorem, stronger than the original solution.
A Diophantine equation is a multivariate polynomial equation over the integers. There are countably infinitely many such equations, and they form a very general language: many difficult problems across number theory and related fields can be equivalently expressed as a single Diophantine equation [13]. One often considers parametric equations of the type $P(a,z_{1},\ldots,z_{\nu})=0$ where $a$ is called the parameter and $z_{1},\ldots,z_{\nu}$ are called unknowns. A Diophantine set $A\subset{\mathbb{N}}$ is then defined to be any set of parameters such that $a\in A$ if and only if there are unknowns $z_{1},\ldots,z_{\nu}\in{\mathbb{Z}}$ with $P(a,z_{1},\ldots,z_{\nu})=0$.
The Davis-Putnam-Robinson-Matiyasevich theorem (1970) [12, Chapter 1] states that every recursively enumerable set is Diophantine; this is a deep relation between number theory and computability theory. An enumeration argument shows that, conversely, every Diophantine set is recursively enumerable. The resulting equivalence implies that there are Diophantine sets which are not decidable, and hence proves that Hilbert’s Tenth Problem is undecidable.
@@ 1.2 Universal Diophantine Equations and Pairs
Using the four-squares substitution for $\nu=9$ over the natural numbers would thus require 36 integer unknowns. A more efficient substitution is possible using the more sophisticated three squares theorem (cf. Section 3.2); this would yield universality for 27 integer unknowns.
@@ 1.3 First Nontrivial Bounds over ${\mathbb{Z}}$
In this article, we formalize the main result leading to the following improvement to only eleven unknowns:
@@
[…]
@@
$(11,1.16258\cdot 10^{63})_{\mathbb{Z}}\qquad\text{and}\qquad(11,9.19593\cdot 10^{53})_{\mathbb{Z}}$
@@
The technique that is used to reduce the number of unknowns to eleven was developed by Sun [19]. However, that work does not consider the degree of the constructed polynomials.
@@
Theorem 1.1 implies the following stronger response to Hilbert’s Tenth Problem.
@@
Consider the class of Diophantine equations with at most 11 unknowns and at most degree $1.63\cdot 10^{63}$. Hilbert’s Tenth Problem is unsolvable for this class: There is no algorithm that can determine for all equations in this class whether they have a solution or not.
Theorem 1.1 crucially rests on the following result. Its proof contains an explicit construction of the polynomial $\tilde{Q}$.
@@
Let $P\in{\mathbb{Z}}[a,z_{1},\ldots,z_{\nu}]$. Then there is an explicit polynomial $\tilde{Q}\in{\mathbb{Z}}[a,z_{1},\ldots,z_{11}]$ such that, for all $a\in{\mathbb{N}}$,
@@
This theorem has been formalized in Isabelle/HOL and is the focus of this paper. It appears, together with Theorem 1.1 and Corollary 1.2, in our recent manuscript [1] which has been written in parallel with the development of its formal proof presented here.
@@ 2 Related Work
Hilbert’s Tenth Problem and related statements have been formalized in a variety of different systems. Note that there are many different but equivalent options to define recursively enumerable sets, either algebraically or more by using some Turing-complete model of computation. In the following, we summarize formalizations of results related to Hilbert’s Tenth Problem, highlighting the different choices authors have made.
In Lean, Carneiro [5] has formalized that exponentiation is Diophantine, which has historically been the last step to be completed in the proof of the DPRM theorem. In Coq, Larchey-Wendling and Forster [9] have formalized the complete DPRM theorem, elegantly using Conway’s FRACTRAN language as a Turing-complete model of computation. They moreover formalize a reduction of their library of unsolvable problems, including two variants of the Halting Problem, to Hilbert’s Tenth Problem. In Mizar, Pak [16] has formalized the DPRM theorem, defining recursively enumerable sets by Davis’ normal form. In Isabelle, Bayer et al. [3] have formalized the DPRM theorem using register machines as a model of computation.
@@
Unfortunately, the published literature often lacked sufficient detail to formalize, and sometimes even to understand, the existing proofs. As such, our detailed mathematical manuscript now reproduces many of the results and proofs cited above. The improved clarity in exposition is a notable product of this formalization.
@@ 3.1 Multivariate Polynomials in Isabelle
$P(z_{0},\ldots,z_{\nu})=\sum_{\begin{subarray}{c}\mathbf{i}\in{\mathbb{N}}^{\nu+1}\\ \|\mathbf{i}\|\leq\delta\end{subarray}}c_{i_{0},\ldots,i_{\nu}}\prod_{s=0}^{\nu}z_{s}^{i_{s}},$ (3)
Although we will not provide typeset proofs of equations (2) and (3), neither equation follows trivially from the definition (1) of the type $\alpha$ mpoly. In Isabelle, we have formalized equation (3) as follows. The term Const C represents a constant polynomial with value $C$, and the term Var s represents the linear polynomial $z_{s}$. The term max_vars P gives the highest variable index that appears in $P$, denoted $\nu$ above.
[…]
4
@@
fixes P :: "$\alpha$ mpoly"
@@
where $f(\mathbf{z})=(f_{0}(\mathbf{z}),\ldots,f_{\nu}(\mathbf{z}))\in{\mathbb{Z}}^{\nu+1}$ is the evaluation of each substitution polynomial at $\mathbf{z}$. In Isabelle, the evaluation morphism is formalized as insertion z P for an assignment of the type z : nat $\Rightarrow\alpha$ and a multivariate polynomial P : $\alpha$ mpoly. Then, equation (4) reads:
@@
lift_definition vars :: "$\alpha$ mpoly $\Rightarrow$ nat set"
is "$\lambda$P. $\bigcup$ (keys keys P)"
definition max_vars :: "$\alpha$ mpoly $\Rightarrow$ nat"
@@
lift_definition total_degree :: "$\alpha$ mpoly $\Rightarrow$ nat"
is "$\lambda$P. Max (insert 0 (($\lambda$m. sum (lookup m) (keys m)) keys P))"
The morphism keys is defined to be the set of all elements of the domain which are mapped to a non-zero value. In particular, keys P is the set of all monomials which appear in $P$ with non-zero coefficient. The morphism lookup denotes function evaluation for finite assignments. Recall also that sum f S $=\sum_{s\in S}f(s)$.
In Isabelle, the Max of a set is not defined for empty sets. Because any constant polynomial has vars (Const c) =absent=$\emptyset$= ∅, inserting the number 00 in the definition of max_vars ensures that the set is always non-empty. When constructing polynomials we always make sure to not skip any variable indices. Moreover, because we work with parametric polynomials, the zeroth variable is always reserved for the parameter $a$. Accordingly, the above definition has the correct semantics for this project.
The definition of total_degree works in analogy to max_vars. Because the zero polynomial has keys (Const 0) =absent=$\emptyset$= ∅, we again include insert 0. However, this formalizes the degree of the zero polynomial as 00. A consequence of this definition is, e.g., that the assumptions $P\neq 0$ and $Q\neq 0$ are now required for the lemma total_degree (P * Q) = total_degree P + total_degree Q.
We formalize appropriate simplification and correctness properties of these definitions for all other functions and morphisms which we use to construct polynomials. This includes exponentiation, finite sums and products, type casting, and substitutions. Because we are concerned with upper bounds of the degree only, it suffices to establish inequalities which require fewer assumptions. Indeed, total_degree (P * Q) $\leq$ total_degree P + total_degree Q is true without any assumptions on $P$ or $Q$.
With respect to the morphism poly_subst introduced above, for example, we formalize the following relative upper bounds. Note that degree P s calculates the degree of $P$ in the variable $z_{s}$ only.
@@
(* Proof omitted *)
@@
4
lemma total_degree_poly_subst:
"total_degree (poly_subst f P)
[…]
(* Proof omitted *)
@@
Recall that Theorem 1.3 asserts that any Diophantine equation can be expressed as a polynomial $\tilde{Q}$ with only eleven integer unknowns. Consequently, the natural choice for the type of $\tilde{Q}$ in Isabelle is int mpoly. Since $\tilde{Q}$ is constructed by substituting various previously defined polynomials into each other, one might expect that these intermediate polynomials also have the type int mpoly.
However, this approach presents notable difficulties when formalizing the Theorems 5.2 and 6.4 which will be presented in the following section. Instead of using meaningful variable names like $a,f,g$, one must work with Var 0, Var 1, Var 2 and frequently refer to the insertion map explicitly. More critically, substituting polynomials into one another requires care to avoid collisions of variable indices. These challenges cannot be fully mitigated through new notation alone.
@@
Using this framework, a polynomial $P$ is initially formalized as a standard Isabelle term of type P :: int => int => ... => int. The command poly_extract P then generates a new term P_poly of type int mpoly and automatically proves a lemma P_correct. This lemma guarantees that evaluating P_poly yields the same result as evaluating the original function P. The extraction succeeds if P is a valid polynomial: constructed from integer constants and variables, addition, multiplication, and any polynomial term C for which a corresponding correctness lemma C_correct exists.
@@ 3.2 The Three Squares Theorem
The three squares theorem states that for every $n\in{\mathbb{N}}$ there exist integers $x,y,z\in{\mathbb{Z}}$ such that $n=x^{2}+y^{2}+z^{2}+z$. It is a consequence of the classical number theory developed in Nathanson’s textbook [15, Lemma 1.9]. Danilkin and Chevalier formalized the three squares theorem as well as its number theory prerequisites in Isabelle [6]. Using this result will constitute the last step in our proof. It serves to ultimately replace the single remaining non-negative integer $n\geq 0$ in Lemma 3.2 below by three true integer unknowns.
@@ 3.3 Relation Combining
$\begin{split}&M_{q}(A_{1},\dots,A_{q},S,T,R,n)\\ &=\prod_{\varepsilon_{1},\dots,\varepsilon_{q}\in\{\pm 1\}^{q}}\left(S^{2}n+T^{2}-S^{2}(2R-1)\left(T^{2}+X_{q}^{q}+\sum_{j=1}^{q}\varepsilon_{j}\sqrt{A_{j}}X_{q}^{j-1}\right)\right)\\ \end{split}$ (5)
@@
[…]
6

### v2

@@ Abstract
In this paper, we contribute a formal verification of this new result. In doing so, we markedly extend the Isabelle AFP entry on multivariate polynomials [17], formalize parts of a number theory textbook [15], and develop classical theory on Diophantine equations [14] in Isabelle. In addition, our work includes metaprogramming infrastructure designed to efficiently handle complex definitions of multivariate polynomials. Our mathematical draft has been formalized while the mathematical research was ongoing, and benefited largely from the help of the theorem prover. We reflect on how the close collaboration between mathematician and computer is an uncommon but promising modus operandi.
@@ 1.1 Hilbert’s Tenth Problem
The tenth problem on David Hilbert’s 1900 list of twenty-three mathematical problems asks for an algorithm to decide the solvability of any arbitrary Diophantine equation [7]. Like many problems on his list, Hilbert’s Tenth Problem has provided inspiration for much mathematical research in the last century, and still does so until today. While the original problem was solved more than fifty years ago [10], this paper presents the formal verification of results related to a new quantitative theorem, stronger than the original solution.
A Diophantine equation is a multivariate polynomial equation over the integers. There are countably infinitely many such equations, and they form a very general language: many difficult problems across number theory and related fields can be equivalently expressed as a single Diophantine equation [13]. One often considers parametric equations of the type $P(a,z_{1},\ldots,z_{\nu})=0$ where $a$ is called the parameter and $z_{1},\ldots,z_{\nu}$ are called unknowns. Then a Diophantine set $A\subset{\mathbb{N}}$ is defined as any set of parameters such that $a\in A$ if and only if there are unknowns $z_{1},\ldots,z_{\nu}\in{\mathbb{Z}}$ with $P(a,z_{1},\ldots,z_{\nu})=0$.
The Davis-Putnam-Robinson-Matiyasevich (DPRM) theorem (1970) [12, Chapter 1] states that every recursively enumerable set is Diophantine; this is a deep relation between number theory and computability theory. An enumeration argument shows that, conversely, every Diophantine set is recursively enumerable. The resulting equivalence implies that there are Diophantine sets which are not decidable, and hence proves that Hilbert’s Tenth Problem is undecidable.
@@ 1.2 Universal Diophantine Equations and Pairs
Using the four-squares substitution for $\nu=9$ over the natural numbers would thus require 36 integer unknowns. A more efficient substitution is possible using the more sophisticated three squares theorem (cf. Section˜3.2); this would yield universality for 27 integer unknowns.
@@ 1.3 First Nontrivial Bounds over $\mathbb{Z}$}
In this article, we formalize the following improvement to only eleven unknowns:
@@
is universal where $\eta(\nu,\delta)=15\,616+233\,856\;\delta+233\,952\;\delta\,(2\delta+1)^{\nu+1}+467\,712\;\delta^{2}\,(2\delta+1)^{\nu+1}$.
@@
$(11,1.68105\cdot 10^{63})\qquad\text{and}\qquad(11,9.50818\cdot 10^{53})$
@@
This theorem has been formalized in Isabelle/HOL and is the focus of this paper. The technique used to reduce the number of unknowns to eleven was developed by Sun [19]. However, that work does not consider the degree of the constructed polynomials.
@@
Theorem˜1.1 implies the following stronger response to Hilbert’s Tenth Problem.
@@
Consider the class of Diophantine equations with at most 11 unknowns and at most degree $1.69\cdot 10^{63}$. Hilbert’s Tenth Problem is unsolvable for this class: There is no algorithm that can determine for all equations in this class whether they have a solution or not.
Theorem 1.1 crucially rests on the following result.
@@
Let $P\in{\mathbb{Z}}[a,z_{1},\ldots,z_{\nu}]$. Then there is an explicit construction of a polynomial $\tilde{Q}\in{\mathbb{Z}}[a,z_{1},\ldots,z_{11}]$ such that, for all $a\in{\mathbb{N}}$,
@@
These theorems and the corollary appear together in our recent manuscript [2] which has been written in parallel to the development of the formal proof presented here.
@@ 2 Related Work
Hilbert’s Tenth Problem and related statements have been formalized in a variety of different systems. Note that there are many different but equivalent options to define recursively enumerable sets, either algebraically or by using some Turing-complete model of computation. In the following, we summarize formalizations of results related to Hilbert’s Tenth Problem, highlighting the different choices that the authors have made.
In Lean, Carneiro [5] has formalized that exponentiation is Diophantine, which has historically been the last step to be completed in the proof of the DPRM theorem. In Coq, Larchey-Wendling and Forster [9] have formalized the complete DPRM theorem, elegantly using Conway’s FRACTRAN language as a Turing-complete model of computation. They moreover formalize a reduction of their library of unsolvable problems, including two variants of the Halting Problem, to Hilbert’s Tenth Problem. In Mizar, Pak [16] has formalized the DPRM theorem, defining recursively enumerable sets by Davis’ normal form. In Isabelle, we have formalized the DPRM theorem using register machines as a model of computation [4].
@@
Unfortunately, the published literature often lacked sufficient detail to formalize, and sometimes even to understand, the existing proofs. As such, our detailed mathematical manuscript [2] now reproduces many of the results and proofs cited above. The improved clarity in exposition is a notable product of this formalization.
@@ 3.1 Multivariate Polynomials in Isabelle
$P(z_{0},\ldots,z_{\nu})=\sum_{\begin{subarray}{c}\mathbf{i}\in{\mathbb{N}}^{\nu+1}\\ \|\mathbf{i}\|\leq\delta\end{subarray}}c_{i_{0},\ldots,i_{\nu}}\prod_{s=0}^{\nu}z_{s}^{i_{s}}.$ (3)
Even if we will not provide proofs of equations (2) and (3), neither equation follows trivially from the definition (1) of the type $\alpha$ mpoly. In Isabelle, we have formalized equation (3) as follows. The term Const C represents a constant polynomial with value $C$, and the term Var s represents the linear polynomial $z_{s}$. The term max_vars P gives the highest variable index that appears in $P$, denoted $\nu$ above.
Listing 1: MPoly_Utils/Poly_Expansions.thy (line 527).
@@
fixes P :: $\alpha$ mpoly"
@@
where $f(\mathbf{z})=(f_{0}(\mathbf{z}),\ldots,f_{\nu}(\mathbf{z}))\in{\mathbb{Z}}^{\nu+1}$ is the evaluation of each substitution polynomial at $\mathbf{z}$. In Isabelle, the evaluation map is formalized as insertion z P for an assignment of the type z : nat $\Rightarrow\alpha$ and a multivariate polynomial P : $\alpha$ mpoly. Then, equation (4) reads:
Listing 3: MPoly_Utils/Substitutions.thy (line 82).
@@
Listing 4: Excerpt from the AFP entry on multivariate polynomials [17].
lift_definition vars :: $\alpha$ mpoly $\Rightarrow$ nat set"
[…]
lift_definition total_degree :: $\alpha$ mpoly $\Rightarrow$ nat"
[…]
Listing 5: MPoly_Utils/Variables.thy (line 185).
definition max_vars :: $\alpha$ mpoly $\Rightarrow$ nat"
@@
The function keys is defined to be the set of all elements of the domain which are mapped to a non-zero value. In particular, keys P is the set of all monomials which appear in $P$ with non-zero coefficient. The map lookup denotes function evaluation for finite assignments. Recall also that sum f S $\sum_{s\in S}f(s)$.
In Isabelle, the Max of a set is not defined for empty sets. Because any constant polynomial has vars (Const C) $\emptyset$, inserting the number 0 in the definition of max_vars ensures that the set is always non-empty. When constructing polynomials we always make sure to not skip any variable indices. Moreover, because we work with parametric polynomials, the zeroth variable is always reserved for the parameter $a$. Accordingly, the above definition has the correct semantics for this project.
The definition of total_degree works in analogy to max_vars. Because the zero polynomial has keys (Const 0) $\emptyset$, we again include insert 0. However, this formalizes the degree of the zero polynomial as 0. A consequence of this definition is, for example, that the assumptions $P\neq 0$ and $Q\neq 0$ are now required for the lemma total_degree (P * Q) = total_degree P + total_degree Q.
We formalize appropriate simplification and correctness properties of these definitions with respect to all other functions that we use to construct polynomials. This includes exponentiation, finite sums and products, type casting, and substitutions. Because we are concerned with upper bounds of the degree only, it suffices to establish inequalities, which require fewer assumptions. Indeed, total_degree (P * Q) $\leq$ total_degree P + total_degree Q is true without any assumptions on $P$ or $Q$.
Remark 3.2.
With the many substitutions of polynomials into polynomials required for our construction, it is useful to clearly state how the calculations of max_vars and total_degree interact with iterated substitutions. For the variables, it is not hard to prove the following, which can be applied recursively.
Listing 6: MPoly_Utils/Substitutions.thy (line 316).
@@
However, for the degree, it becomes necessary to fully expand all substitutions to obtain tight bounds. For example, consider first evaluating the degree of $P=Q_{1}Q_{2}+Q_{3}$. Given an arbitrary substitution $f$, this polynomial can have many different degrees in the variables $z_{s}$. [^1] Consider, for example, the two substitutions $f=(Q_{1},Q_{2},Q_{3})=(z_{1},z_{2},z_{2}^{2})$ and $f=(z_{1}^{5},z_{2}^{3},z_{2})$. Of course, there is an explicit formula for the total degree of this $P$. However, for both arbitrary $f$ and arbitrary $P$, an elementary recursive step, analogous to the above lemma for variables, cannot suffice.
Instead, we rely on the following auxiliary definition, which will be used by the automation tools introduced below. We abbreviate Max (insert 0 -) by just Max0 -.
Listing 7: MPoly_Utils/Total_Degree_Env.thy (line 12).
lift_definition total_degree_env :: "(nat $\Rightarrow$ nat) $\Rightarrow$ $\alpha$ mpoly $\Rightarrow$ nat"
@@
Here, $e:{\mathbb{N}}\rightarrow{\mathbb{N}}$ denotes an “environment” of degrees $\delta_{s}$ for each basic variable $z_{s}$. Note the special case total_degree_env ($\lambda$_. 1) P = total_degree P. The environment is updated with the degree of each $f_{s}$ whenever we encounter poly_subst f P, modeling our algorithm to compute the degree of an arbitrary multivariate polynomial.
@@
Recall that Theorems˜1.1 and 1.3 assert that any Diophantine equation can be expressed as a polynomial $\tilde{Q}$ with only eleven integer unknowns. Consequently, the natural choice for the type of $\tilde{Q}$ in Isabelle is int mpoly. Since $\tilde{Q}$ is constructed by substituting various previously defined polynomials into each other, one might expect that these intermediate polynomials also have the type int mpoly.
However, this approach presents notable difficulties when formalizing the Theorems˜5.2 and 6.4 which will be presented in the following section. Instead of using meaningful variable names like $a,f,g$, one must work with Var 0, Var 1, Var 2 and frequently refer to the insertion map explicitly. More critically, substituting polynomials into one another requires care to avoid collisions of variable indices. These challenges cannot be fully mitigated through new notation alone.
@@
Using this framework, a polynomial $P$ is initially formalized as a standard Isabelle term of type P :: int => int => ... => int. The command poly_extract P, implemented by Anna Danilkin in our workgroup, then generates a new term P_poly of type int mpoly and automatically proves a lemma P_correct. This lemma guarantees that evaluating P_poly yields the same result as evaluating the original function P. The extraction succeeds if P is a valid polynomial: constructed from integer constants and variables, addition, multiplication, exponentiation with a natural number, and any term for which appropriate translations to a polynomial already exist from previous applications of poly_extract.
Calculating the total degree of the final universal Diophantine equation $\tilde{Q}$ and proving its correctness also posed significant challenges due to the depth of iterated substitutions. The correctness proofs, in particular, require thousands of invocations of the various simplification lemmas $\deg(PQ)\leq\deg P+\deg Q$ for multiplication, $\deg(P+Q)\leq\max\{\deg P,\deg Q\}$ for addition, and $\deg P^{n}\leq n\deg P$ – specifically, one invocation for each arithmetic operation that appears in our construction. To automate both the calculation of the degree and an accompanying proof of correctness, Annie Yao developed the command poly_degree.
When invoked on a term of type P_poly :: int mpoly, this command recursively computes (an upper bound for) the degree of this term, called P_poly_degree, and automatically proves a lemma P_poly_degree_correct. The lemma states that the total degree of P_poly is indeed bounded above by P_poly_degree. The calculation and proof succeed when P_poly is constructed from basic arithmetic operators, variables, constants (including arbitrary Isabelle terms of type int) and, crucially, polynomial substitutions.
@@ 3.2 The Three Squares Theorem
The three squares theorem states that for every $n\in{\mathbb{N}}$ there exist integers $x,y,z\in{\mathbb{Z}}$ such that $n=x^{2}+y^{2}+z^{2}+z$. It is a consequence of the classical number theory developed in Nathanson’s textbook [15, Lemma 1.9]. Danilkin and Chevalier [6] formalized the three squares theorem as well as its number theory prerequisites in our workgroup. Using this result will constitute the last step in our proof. It serves to ultimately replace the single remaining non-negative integer $n\geq 0$ in Lemma 3.3 below by three true integer unknowns.
@@ 3.3 Relation Combining
[…]
@@
The shorthand notation in equation (5) defines a polynomial because the square roots cancel out in the product (we omit a proof here). In our case, we do not require the general form of this result and instead only its specialization to $q=3$. Using the automation built into poly_extract as discussed above, it moreover suffices to simply spell out the explicit function $\texttt{M3}:{\mathbb{Z}}^{7}\rightarrow{\mathbb{Z}}$. Lemma 3.3 is then formalized as follows.
Listing 8: Nine_Unknowns_N_Z/Matiyasevich_Polynomial.thy (line 205).